40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 283     Entry time: Mon Aug 16 16:36:35 2010
Author: Frank 
Type: Summary 
Category: Laser 
Subject: laser, PMC, refcav re-characterization and recent changes 

laser & PMC:

re-characterized the laser and PMC

new laser temp setpoint : 44.1 C  - almost centered between mode hops,
also PMC resonance close to that with centered PZT offset

  • PMC PZT OFFSET (PMC_RAMP) = -2.7V = 145V @ PZT (300V max)
  • PMC PZT tuning: 2.24V / FSR (714MHz)  for EXT DC INPUT
  • PMC TRANSPD close to limit: changed gain, now 3.32V=24.6mW  (7.41mW/V)

current values for TEM00:
RCAV resonant at 0.0182V
ACAV resonant at 0.0566V

ACAV:

  • re-aligned and re-modematched ACAV as old mode matching was only 47%
  • now DC voltage in reflection is 359mV if not locked and 42mV if locked. -> now 88% coupled into cavity
  • position of lenses very sensitive -> changing position a millimeter makes huge difference but no space for different mode matching
  • changed position of RF photodiode a little bit
  • beam size on PD now a little bit smaller
  • beam dump added for reflected beam from RF photodiode
  • changed setpt temp to 36.5 C, since the box changed the difference between both cavities. So the actual temp read and stabilized is not the (only) one seen by the cavity. The gradient seems to be important too.
ELOG V3.1.3-